OpenStudy (anonymous):

Please Help!!! The life times, Y in years of lightbulbs follow an exponential distribution with a mean of 0.7 . A tester makes random observations of the life times of this lightbulbs and records them one by one as either a success if the life time exceeds 1 year, or as a failure otherwise. Find probability of second success occurring in the 8th observation given that first success occurred in the 3rd observation. Find the probability that the first success occurs in an odd-numbered observation. That is, the success occurs in the 1st or 3rd or 5th or 7th (and so oon) observation

OpenStudy (anonymous):

Since \(Y\) follows an exponential distribution, the corresponding density function \[f(y)=\begin{cases}\dfrac{1}{\beta}e^{-y/\beta}&\text{for }y\ge0\\0&\text{otherwise}\end{cases}\] and the corresponding distribution function is \[F(y)=\begin{cases}1-e^{-y/\beta}&\text{for }y\ge0\\0&\text{otherwise}\end{cases}\] where \(\beta\) is the mean. The observations will have geometric probability attributed to them, so first you need to the find the success probability, \(p\), which is equal to \(P(Y>1)\). \[\begin{align*}p&=P(Y>1)\\&=1-P(Y<1)\\ &=1-F(1)\\ &=1-e^{-1/0.7}\\ &\approx 0.76 \end{align*}\] The setup for the first probability is explained pretty well in this link: http://math.stackexchange.com/a/134220 For the second probability, if we denote \(A_n\) to mean "a success occurs on the \(n\) observation," then, since the desired observations are mutually exclusive, \[\begin{align*}P\left(\bigcup_{k=0}^{\infty}A_{2k+1}\right)&=\sum_{k=0}^\infty P(A_{2k+1})\\ P(A_1\cup A_3\cup\cdots)&=P(A_1)+P(A_3)+\cdots\end{align*}\] Since \(A_n\) follows a geometric distribution, you have \[\sum_{k=0}^\infty P(A_{2k+1})=\sum_{k=0}^\infty(1-p)^{2k+1-1}p\] DISCLAIMER: I'm not totally sure about the set up for the second part, but it somewhat makes sense to me.

OpenStudy (anonymous):

Oh I just noticed, the \(p\) I found is incorrect. It should be \[p=1-\left(1-e^{-1/.7}\right)\approx 0.24\]